OpenStudy (anonymous):

Please help, I'll give you a medal! From a committee of six girls and four boys, a name is selected. Then another name is selected. What is the probability that both people drawn from will be girls. A.9/25 B.2/15 C.3/5 D.1/3

OpenStudy (y1066380):

Total number of boys and girls=6+4=10 P(2 girls) =(6/10)*(5/9) =1/3 =D Do you know why?

OpenStudy (anonymous):

When you draw the first name, its probability of being a girl's name is 6/10, since there are 6 girls and a total of 10 people After selecting the first girl, you are left with the names of 5 girls and a total of 9 people. Hence when you go on selecting the second name, its probability of being a girl becomes 5/9 The probability of both the above events occurring together is 6/10 * 5/9 = 1/3